You are trying to get an idea of how predictable a firm future FCF’s will be so you look at the last 10 years FCF’s and find that the average was 25 million and that the standard deviation was 18 million. You decide to find the range of outcomes 95% of the time. How should you interpret this?

a. The range of outcomes are from 7 to 43 million

b. The range of outcomes are from -11 to 61 million and the smaller the standard deviation the lower the total risk

c. The range of outcomes are from -11 to 61 million and the smaller the standard deviation the higher the total risk

d. The range of outcomes are from 5 to 45 million
Answer #1
Mean FCF 25 million
Standard Deviation of FCF 18 million
As per 68-95-99.7 rule
95% probability will be 2 sigma events
Z=(X-Mean)/Standard Deviation
Value of Z with a range of -2 and +2 will have probability of 95%
Range for 95% outcomes:
Lower Range :(X-25)=-2*18; X=-11
Higher Range :(X-25)=2*18: X=61
Hence, the range of outcome are from -11 to 61
Standard deviation indicates the amount of risk
Higher the standard deviation , higher the risk
Hence, Lower the Standard Deviation , Lowr the risk
ANSWER :
b. The range of outcomes are from -11 to 61 million
and the smaller the standard deviation the lower the total risk